# equal contribution

### v1

@@ Abstract
Recent techniques for text-to-4D generation synthesize dynamic 3D scenes using supervision from pre-trained text-to-video models. However, existing representations for motion, such as deformation models or time-dependent neural representations, are limited in the amount of motion they can generate—they cannot synthesize motion extending far beyond the bounding box used for volume rendering. The lack of a more flexible motion model contributes to the gap in realism between 4D generation methods and recent, near-photorealistic video generation models. Here, we propose TC4D: trajectory-conditioned text-to-4D generation, which factors motion into global and local components. We represent the global motion of a scene’s bounding box using rigid transformation along a trajectory parameterized by a spline. We learn local deformations that conform to the global trajectory using supervision from a text-to-video model. Our approach enables the synthesis of scenes animated along arbitrary trajectories, compositional scene generation, and significant improvements to the realism and amount of generated motion, which we evaluate qualitatively and through a user study. Video results can be viewed on our website: https://sherwinbahmani.github.io/tc4d.
@@ 1 Introduction
Recent advances in video generation models [11, 93, 9, 77, 33, 10, 95, 99] enable the synthesis of dynamic visual content with unprecedented fidelity. These advances have promoted the rapid development of techniques for 3D video generation (see Fig. 1), which use supervision from video generation models to produce animated 3D scenes, often with input conditions, e.g., video [39], image [71], or text prompt [78]. With applications including visual effects, virtual reality, and industrial design, realistic and controllable generation of 4D content has broad potential impact.
@@
Although methods for 4D generation show immense promise, they are yet limited in the quality and realism of synthesized motion effects. Specifically, these methods animate a volumetric scene representation via learned deformation [50, 71, 108] or appearance changes predicted by a neural representation [115, 116, 39, 4] or generative model [63]. Since the scene is confined to a particular region—e.g., a 3D bounding box containing the volumetric representation—the synthesized motion is highly local, and the magnitude of motion is typically much smaller than the scale of the scene itself. For example, existing techniques can synthesize characters that walk in place, gesticulate, or sway, but they cannot move around a scene or exhibit global motion effects in general. Overall, there is a stark contrast in realism between motion effects synthesized with video generation models and existing techniques for 4D generation (see Fig. 2).
@@
Figure 2: Comparing text-to-video and text-to-4D generation. Existing methods for video generation (e.g., VideoCrafter2 [14]) create realistic objects with motion at both global and local scales. For example, a generated unicorn or corgi moves across the camera field of view (global motion) in concert with the animation of its legs and body (local motion). However, existing text-to-4D methods [4] only generate a limited amount of motion on the spot and do not learn global motion in 3D space.
@@ 2 Related Work
This section reviews the most relevant areas of our work: text-to-video, text-to-3D, and text-to-4D generative models. For a comprehensive overview of recent advances in diffusion models and 4D generation, we refer readers to [66] and [111].
@@
Our work builds on the advancements in 2D image and video generative models, particularly focusing on text-to-video generation and its challenges, as highlighted by recent literature [8, 66]. Earlier works favored GANs [79, 2], but the trend has shifted towards diffusion models for their superior performance in video generation [11, 54]. However, most existing text-to-video models still exhibit a significant quality gap with text-to-image models due to dataset and computational constraints. Hybrid training approaches combining image and video datasets have been explored to overcome these limitations, employing techniques for enhancing spatial and temporal resolution through pixel-space and latent space upsampling [7, 95, 104, 33, 28, 32, 96, 117]. The emergence of transformer-based autoregressive models signifies a leap forward, enabling the creation of videos of arbitrary lengths [86, 11, 57, 59]. Recent efforts have also focused on fine-tuning text-to-image models for video data and decoupling content and motion generation for enhanced quality and flexibility [10, 77, 99, 28, 9]. Additionally, research aimed at increasing video generation controllability has shown promising results, allowing for explicit manipulation of various video aspects [98, 92, 75, 106, 6].
Despite advancements, high-quality video synthesis remains challenging, with a notable quality disparity between open-source models and top-performing proprietary ones [1, 93]. Although the recent open-source SVD [9] can synthesize high-quality videos from image through meticulous data curation, we use VideoCrafter2 [14] for our experiments due to its support for text conditioning.
@@
Early works on 3D generation leveraged the development of GAN algorithms to create realistic 3D objects of single categories, e.g., human or animal faces, cars, etc. These 3D GAN approaches [19, 12, 62, 74, 3], however, were limited in the class of scenes they could generate and thus were not suitable for text-based generation. Initial attempts at data-driven, text-conditioned methods used CLIP-based supervision [70], which enabled synthesis or editing of 3D assets [15, 36, 73, 37, 25, 90]. These techniques evolved into the most recent approaches, which optimize a mesh or radiance field based on Score Distillation Sampling (SDS) and its variants using pre-trained diffusion models [67, 97, 48, 16, 47, 91, 46, 31, 107, 53, 110, 40, 43, 80]. These SDS-based methods are improving, generating increasingly high-quality 3D assets. The quality of generated 3D structures has been further improved by applying diffusion models that consider multiple viewpoints [49, 52, 76, 21, 51, 41, 88, 34, 81]. These models alleviate the so-called ”Janus” problem, where objects are generated with extra heads, limbs, etc. Alternatively, recent advancements have seen a shift towards using diffusion or transformer models to lift an input 2D image into a 3D representation for novel-view synthesis [13, 83, 26, 55, 109, 84, 69, 56, 89]. Other recent works [35, 44, 103, 102, 113, 30, 38, 101, 82, 85, 68] consider feed-forward 3D generation methods by first generating novel view images and converting them into 3D using transformer architectures. Another line of work models compositional 3D scenes [65, 18, 5, 17, 87, 20, 118, 23, 114] by decomposing the scene into objects. While producing impressive results, these techniques do not support generating 4D scenes, i.e., their generated objects/scenes are static. We adapt MVDream [76] and ProlificDreamer [97] to obtain an initial high-quality and static 3D model, which we animate along a trajectory using supervision from a video diffusion model.
@@
Our research aligns closely with SDS-based 4D generation methodologies that are initialized with a 3D model and subsequently animate it using text-to-video supervision. Within the last year, significant progress from initial results in 4D generation [78] has been made by a slew of recent methods [71, 50, 4, 116, 24, 105, 39, 105]. These techniques have notably enhanced both geometry and appearance by integrating supervision from various text-to-image models. More recent endeavors [71, 115, 108, 63, 116, 50, 24, 112] have employed similar SDS-based strategies for image or video-conditioned 4D generation. Still, all of these approaches yield 4D scenes that remain stationary at a single location and are limited to short and small motions, such as swaying or turning. Our objective, however, is to generate large-scale motions that more accurately mimic real-world movements.
@@ 3.3 Modeling Local Motion with Trajectory-Aware VSDS
We capture local motion by optimizing the deformation model with guidance from a trajectory-aware verison of VSDS using a text-to-video diffusion model. To regularize the generated deformation field we use a smoothness penalty, and we introduce a motion annealing technique, which empirically improves realism and temporal consistency.
@@
One challenge with using VSDS to generate dynamic objects along a trajectory is that the pretrained video diffusion models generate a finite number of $N$ frames and thus can only represent limited range of motion. Naively sampling $N$ timesteps across a long trajectory would require the diffusion model to represent long-range motion, different from its training data.
@@ 3.4 Implementation details
Additionally, we find that VSDS gradients for diffusion timesteps where $t_{\text{d}}\approx 1$ introduce jittery, high-frequency motion into the deformation field. We address this issue by annealing the sampling interval of the diffusion timesteps during the course of optimization, which improves the temporal coherence of local motion. At the start of training we sample compute $\nabla\mathcal{L}_{\text{4D}}$ by uniformly sampling $t_{d}\in[0.02,0.98]$. This interval is linearly decayed during training to $t_{d}\in[0.02,0.5]$.
@@ 4 Experiments
Motion quality is extremely difficult to visualize. In this manuscript, we focus on presenting a detailed analysis from our quantitative evaluation and include only the most representative visual examples. We highly encourage the readers to check the videos in the supplemental material for a conclusive quality assessment.
@@ 4.1 Metrics
Since there is no automated metric for 4D generation, we evaluate TC4D with a user study. The method is compared quantitatively and qualitatively to a modified version of 4D-fy [4]. That is, we animate the output of 4D-fy by rigidly transforming it to follow the same trajectory as used for TC4D. In total we provide comparisons using 33 scenes generated from different text prompts, and we conduct an extensive ablation study on a subset of these text prompts. We note that text-to-4D generation remains a computationally expensive procedure; our evaluation uses a similar number of text prompts as previous work (e.g., 28 [78], or 17 [4]).
@@
We recruit a total of 20 human evaluators and show the evaluators two video clips generated using the same text prompt with TC4D and 4D-fy. We ask the evaluators to specify their overall preference and their preference in terms of appearance quality (AQ), 3D structure quality (SQ), motion quality (MQ), motion amount (MA), text alignment (TA) (see the supplement for the detailed protocol for the study). The results of the user study are shown in Table 1 and we include further details in the supplement. Each row of the table reports the percentage of evaluations which ranked T4CD over 4D-fy or the ablated approach, i.e., a percentage above 50% indicates T4CD is preferred. Statistically significant results ($p<0.05$) are indicated based on a $\chi^{2}$ test.
@@ 4.2 Results
We visualize the generated 4D scenes by rendering and compositing images along the trajectory from multiple camera viewpoints in Fig. 4. TC4D shows significantly more motion and also more realistic motion than 4D-fy—this is especially noticeable for scenes in which TC4D synthesizes walking animations that are coherent with the trajectory, but 4D-fy produces only a small amount of relatively incoherent motion. In the user study, participants indicated a statistically significant preference for results from TC4D for all metrics except structure quality. In terms of overall preference, 85% of comparisons were in favor of T4CD. We refer the reader to the video clips included in the supplement, which provide a clear visualization of the advantages of the proposed approach.
@@ 4.3 Ablation Study
Alternatively, entirely removing the trajectory reduces the method to the conventional text-to-4D generation setting with deformation-based motion. That is, we have $\mathcal{N}_{\text{4D}}=\mathcal{N}_{\text{3D}}(\mathcal{D}(\mathbf{x},t))$. This experiment is indicated in Table 1 as “w/o local trajectory”. While the method still learns to incorporate some motion without the trajectory, we find that the trajectory is key to producing more complex animations such as walking. This leads to a clear advantage of our method in the user study not only for metrics related to motion (MQ and MA), but also affect the text alignment (TA) and overall preference (last column).
@@
We use trajectory-aware VSDS to ensure that all regions of the trajectory receive adequate supervision from the text-to-video model. Specifically, this procedure involves breaking the trajectory down into segments, and then sampling one segment for which we render the video clips used to compute the VSDS gradients. We ablate this component by rendering video frames from a specific camera viewpoint sampled at equal steps $t$ along the entire trajectory. Then the VSDS gradients are computed using the resulting video. Overall, we find this approach results in too coarse a sampling of the trajectory and fails to synthesize smooth motion. This effect can be most clearly seen in the video clips included in the supplement. In Table 1, this method is indicated as “w/o traj.-aware VSDS”. Most users are in favor of our method for higher motion quality.
@@
We assess the impact of removing the smoothness penalty from the deformation field (Eq. 6). This regularization term, proposed by Zheng et al. [116], helps to dampen high-frequency displacements in the deformation field and leads to smoother motion. As the row “w/o smooth penalty” shows in Table 1, the responses from user study participants indicate that removing this penalty has only a modest negative effect, especially compared to the timestep annealing ablation, which we discuss next.
@@
We demonstrate this approach on a scene consisting of a column of fire that is generated with the text prompt, “a flame getting larger” (Fig. 5). Generating this scene without the trajectory and without bounding box scaling results in a flame that has little variation in size. Alternatively, we animate the flame along an upwards trajectory, but this approach simply translates the generated scene rather than changing the size. Combining translation and scaling in the vertical dimension allows us to keep the center of the bounding box at a consistent location, and with this technique, the generated flame appears to grow in size. Overall, adding the scaling term offers additional flexibility in the types of scenes that can be generated with trajectory conditioning.
@@ 5 Conclusion
Overall our work takes important steps in making motion synthesis for text-to-4D models more realistic and expressive. Our framework enables scene-scale motion of entities within compositional 4D scenes with trajectory conditioning, achieving meaningful improvement over prior works that show motions at fixed locations. The proposed factorized motion model opens several avenues for future work. In particular, we envision extensions to generate scenes with multiple interacting objects, adding motion to existing in-the-wild 3D assets, and synthesis of motion across unbounded, large-scale scenes. Our current approach assumes that the initial layout and trajectory are given, but it may also be possible to optimize or generate the layouts and trajectories in an end-to-end fashion. An ultimate holistic approach would include an end-to-end pipeline for synthesizing both initial objects and their trajectories for compositional 4D scene generation, which we continue to explore. Finally, designing automated text-to-4D metrics by temporally extending text-to-3D metrics [100] is an important direction for future work.
@@
Our approach can automatically generate realistic, animated 3D scenes; such techniques can be misused to promulgate realistic fake content for the purposes of misinformation. We condemn misuse of generative models in this fashion and emphasize the importance of research to thwart such efforts (see, e.g., Masood et al. [58] for an extended discussion of this topic). It is to note, however, that the current level of 4D generation technology still allows humans to easily identify the generated content. In the future, we expect quality improvements such that generated 4D scenes could become indistinguishable from real content. The literature on fake content detection [29, 22, 94] claims that it is easy for now to spot generated content. We fully support the development of algorithms for detection of generated content to help prevent potential misuse of generative AI technologies.
6 Acknowledgements

### v3

@@ Abstract
Recent techniques for text-to-4D generation synthesize dynamic 3D scenes using supervision from pre-trained text-to-video models. However, existing representations, such as deformation models or time-dependent neural representations, are limited in the amount of motion they can generate—they cannot synthesize motion extending far beyond the bounding box used for volume rendering. The lack of a more flexible motion model contributes to the gap in realism between 4D generation methods and recent, near-photorealistic video generation models. Here, we propose TC4D: trajectory-conditioned text-to-4D generation, an approach that factors motion into global and local components. We represent the global motion of a scene’s bounding box using rigid transformation along a trajectory parameterized by a spline. We learn local deformations that conform to the global trajectory using supervision from a text-to-video model. Our approach enables synthesis of scenes animated along arbitrary trajectories, compositional scene generation, and significant improvements to the realism and amount of generated motion, which we evaluate qualitatively and through a user study. Video results can be viewed on our website: https://sherwinbahmani.github.io/tc4d.
@@ 1 Introduction
Recent advances in video generation models [11, 90, 9, 74, 31, 10, 91, 95] enable the synthesis of dynamic visual content with unprecedented fidelity. These advances have prompted the rapid development of techniques for 3D video generation (see Fig. 1), which use supervision from video generation models to produce animated 3D scenes, conditioned on an input video [37], image [68], or text prompt [75]. With applications including visual effects, virtual reality, and industrial design, realistic and controllable generation of 4D content has broad potential impact.
@@
Although methods for 4D generation show immense promise, they are yet limited in the quality and realism of synthesized motion effects. Specifically, these methods animate a volumetric scene representation via learned deformation [48, 68, 104] or appearance changes predicted by a neural representation [111, 112, 37, 4] or generative model [60]. Since the scene is confined to a particular region—e.g., a 3D bounding box containing the volumetric representation—the synthesized motion is highly local, and the magnitude of motion is typically much smaller than the scale of the scene itself. For example, existing techniques can synthesize characters that walk in place, gesticulate, or sway, but synthesized outputs cannot move around a scene or exhibit global motion effects in general. Overall, there is a stark contrast in realism between motion effects created with video generation models and existing techniques for 4D generation (see Fig. 2).
@@
Figure 2: Comparing text-to-video and text-to-4D generation. Existing methods for video generation (e.g., VideoCrafter2 [14]) create realistic objects with motion at both global and local scales. For example, a generated unicorn or corgi moves across the camera field of view (global motion) in concert with the animation of its legs and body (local motion). However, existing text-to-4D methods [4] only generate a limited amount of local motion and do not learn global motion in 3D space.
@@ 2 Related Work
Our work is most directly connected to text-to-video, text-to-3D, and text-to-4D generative models. For a comprehensive overview of recent advances in diffusion models and 4D generation, we refer readers to [63] and [107].
@@
We build on recent advances in 2D image and video generative models, with methods for text-to-video generation being especially relevant to our approach [8, 63]. While earlier methods for image and video generation favor GANs [76, 2], recent work shifts towards using diffusion models for their superior performance in video generation [11]. Still, most existing text-to-video models exhibit a significant quality gap compared to text-to-image models due to dataset and computational constraints. Hybrid training approaches combining image and video datasets have been explored to overcome these limitations, employing techniques for enhancing spatial and temporal resolution through pixel-space and latent space upsampling [7, 91, 100, 31, 27, 30, 92, 113]. The emergence of transformer-based autoregressive models provides a leap forward, enabling the creation of videos of arbitrary lengths [83, 11, 54, 56]. Recent efforts have also focused on fine-tuning text-to-image models for video data and decoupling content and motion generation for enhanced quality and flexibility [10, 74, 95, 27, 9]. Additionally, research aimed at increasing video generation controllability has shown promising results, allowing certain properties of an output video to be explicitly manipulated [94, 89, 72, 102, 6].
Despite recent advances, high-quality video synthesis remains challenging, with a notable quality disparity between open-source models and top-performing proprietary ones [1, 90]. Although the recent open-source method Stable Video Diffusion [9] can synthesize high-quality videos from an input image, we use VideoCrafter2 [14] for our experiments due to its support for text conditioning.
@@
Early work on 3D generation leveraged the development of GAN algorithms to create realistic 3D objects of single categories, e.g., human or animal faces, cars, etc. These 3D GAN approaches [19, 12, 59, 71, 3], however, were limited in the class of scenes they could generate and thus were not suitable for text-based generation. Initial attempts at data-driven, text-conditioned methods used CLIP-based supervision [67], which enabled synthesis or editing of 3D assets [15, 34, 70, 35, 24, 87]. These techniques evolved into the most recent approaches, which optimize a mesh or radiance field based on Score Distillation Sampling (SDS) and its variants using pre-trained diffusion models [64, 93, 46, 16, 45, 88, 44, 29, 103, 51, 106, 38, 41, 77]. These SDS-based methods are improving, generating increasingly high-quality 3D assets. The quality of generated 3D structures has been further improved by applying diffusion models that consider multiple viewpoints [47, 50, 73, 21, 49, 39, 85, 32, 78]. These models alleviate the so-called ”Janus” problem, where objects are generated with extra heads, limbs, etc. Recent methods have shifted toward using diffusion or transformer models to lift an input 2D image into a 3D representation for novel-view synthesis [13, 80, 25, 52, 105, 81, 66, 53, 86]. Other recent work [33, 42, 99, 98, 109, 28, 36, 97, 79, 82, 65] considers feed-forward 3D generation methods by first generating novel view images and converting them into 3D using transformer architectures. Another line of work models compositional 3D scenes [62, 18, 5, 17, 84, 20, 114, 22, 110] by decomposing the scene into objects. While these techniques for 3D generation produce impressive results, they do not support generating 4D scenes, i.e., their generated objects/scenes are static. We adapt MVDream [73] and ProlificDreamer [93] to obtain an initial high-quality and static 3D model, which we animate along a trajectory using supervision from a video diffusion model.
@@
Our research aligns closely with SDS-based 4D generation methodologies that are initialized with a 3D model and subsequently animated using text-to-video supervision. Within the last year, significant progress from initial results in 4D generation [75] has been made by a slew of recent methods [68, 48, 4, 112, 23, 101, 37, 101]. These techniques have notably enhanced both geometry and appearance by integrating supervision from various text-to-image models. More recent endeavors [68, 111, 104, 60, 112, 48, 23, 108] have employed similar SDS-based strategies for image or video-conditioned 4D generation. Still, all of these approaches yield 4D scenes that remain stationary at a single location and are limited to short and small motions, such as swaying or turning. Our objective, however, is to generate large-scale motions that more accurately mimic real-world movements.
@@ 3.3 Modeling Local Motion with Trajectory-Aware VSDS
We capture local motion by optimizing the deformation model with guidance from a trajectory-aware version of VSDS using a text-to-video diffusion model. To regularize the generated deformation field we use a smoothness penalty, and we introduce a motion annealing technique, which empirically improves realism and temporal consistency.
@@
One challenge with using VSDS to generate dynamic objects along a trajectory is that the pre-trained video diffusion models generate a finite number of $N$ frames and thus can only represent limited range of motion. Naively sampling $N$ timesteps across a long trajectory would require the diffusion model to represent long-range motion, different from its training data.
@@ 3.4 Implementation details
Additionally, we find that VSDS gradients for diffusion timesteps where $t_{\text{d}}\approx 1$ introduce jittery, high-frequency motion into the deformation field. We address this issue by annealing the sampling interval of the diffusion timesteps during the course of optimization, which improves the temporal coherence of local motion. At the start of training we compute $\nabla\mathcal{L}_{\text{4D}}$ by uniformly sampling $t_{d}\in[0.02,0.98]$. This interval is linearly decayed during training to $t_{d}\in[0.02,0.5]$.
@@ 4 Experiments
Motion quality is extremely difficult to visualize. In this manuscript, we focus on presenting a detailed analysis of our quantitative evaluation and include only the most representative visual examples. We highly encourage the readers to check the videos in the supplemental material for a conclusive quality assessment.
@@ 4.1 Metrics
Since there is no automated metric for 4D generation, we evaluate TC4D with a user study. We refer to Sec. 3 of the appendix for additional CLIP score evaluations. The method is compared quantitatively and qualitatively to a modified version of 4D-fy [4]. That is, we animate the output of 4D-fy by rigidly transforming it to follow the same trajectory as used for TC4D. We further compare to the concurrent work DreamGaussian4D [68] in Sec. 3 of the appendix. In total we provide comparisons using 33 scenes generated from different text prompts, and we conduct an extensive ablation study on a subset of these text prompts. Our evaluation uses a similar number of text prompts as previous work (e.g., 28 [75], or 17 [4]).
@@
We recruit a total of 20 human evaluators and show the evaluators two video clips generated using the same text prompt with TC4D and 4D-fy. We ask the evaluators to specify their overall preference and their preference in terms of appearance quality (AQ), 3D structure quality (SQ), motion quality (MQ), motion amount (MA), text alignment (TA) (see the supplement for the detailed protocol for the study). The results of the user study are shown in Table 1 and we include further details in the supplement. Each row of the table reports the percentage of evaluations that ranked T4CD over 4D-fy or the ablated approach, i.e., a percentage above 50% indicates T4CD is preferred. Statistically significant results ($p<0.05$) are indicated based on a $\chi^{2}$ test.
@@ 4.2 Results
We visualize the generated 4D scenes by rendering and compositing images along the trajectory from multiple camera viewpoints in Fig. 4. TC4D shows significantly more motion and also more realistic motion than 4D-fy—this is especially noticeable for scenes in which TC4D synthesizes walking animations that are coherent with the trajectory, but 4D-fy produces only a small amount of relatively incoherent motion. In the user study, participants indicated a statistically significant preference for results from TC4D for all metrics except structure quality. In terms of overall preference, 85% of comparisons were in favor of T4CD. We refer the reader to the video clips included in the supplement, which provide a clear visualization of the advantages of the proposed approach. In addition to these results based on user-defined trajectories, we show results for automated end-to-end generation using large language models in Sec. 4 of the appendix. Moreover, we provide geometry visualizations in Sec. 5 of the appendix
@@ 4.3 Ablation Study
Alternatively, entirely removing the trajectory reduces the method to the conventional text-to-4D generation setting with deformation-based motion. That is, we have $\mathcal{N}_{\text{4D}}=\mathcal{N}_{\text{3D}}(\mathcal{D}(\mathbf{x},t))$. This experiment is indicated in Table 1 as “w/o trajectory”. While the method still learns to incorporate some motion without the trajectory, we find that the trajectory is key to producing more complex animations such as walking.
@@
We use trajectory-aware VSDS to ensure that all regions of the trajectory receive adequate supervision from the text-to-video model. Specifically, this procedure involves breaking the trajectory down into segments, and then sampling one segment for which we render the video clips used to compute the VSDS gradients. Then the VSDS gradients are computed using the resulting video. Overall, we find this approach results in too coarse a sampling of the trajectory and fails to synthesize smooth motion. In Table 1, this method is indicated as “w/o traj.-aware VSDS”.
Global transformations.
To demonstrate that VSDS is trajectory dependent, we conduct an ablation where we remove global transformations during training, but still sample the time dimension in a segmented way as in trajectory-aware VSDS. The local deformations in this approach are significantly reduced.
@@
We assess the impact of removing the smoothness penalty from the deformation field (Eq. 6). This regularization term, proposed by Zheng et al. [112], helps to dampen high-frequency displacements in the deformation field and leads to smoother motion. As the row “w/o smooth penalty” shows in Table 1, the responses from user study participants indicate that removing this penalty has only a modest negative effect.
@@
We demonstrate this approach on a scene consisting of a column of fire that is generated with the text prompt, “a flame getting larger” (Fig. 5). Generating this scene without the trajectory and without bounding box scaling results in a flame that has little variation in size. Alternatively, we animate the flame along an upwards trajectory, but this approach simply translates the generated scene rather than changing the size. Combining translation and scaling in the vertical dimension allows us to keep the center of the bounding box at a consistent location, and with this technique, the generated flame appears to grow in size.
@@ 5 Conclusion
Overall our work takes important steps in making motion synthesis for text-to-4D models more realistic and expressive. Our framework enables scene-scale motion of entities within compositional 4D scenes with trajectory conditioning, achieving improvement over prior work that shows motion at fixed locations.
The main limitations of our results are imperfect leg motion, foot sliding, or asymmetric and not fully cyclic motion. The motion imperfections stem from the quality of publicly available pre-trained text-to-video models used for VSDS. Also, VideoCrafter2 outputs 16-frame videos that may be too short to capture fully cyclic motion (e.g., a complete walking cycle), contributing to this limitation. We expect that TC4D will benefit from rapid improvements in 2D video generation, and we open source our code to spur follow-on work.
The proposed factorized motion model opens several avenues for future work. In particular, we envision extensions to generate scenes with multiple interacting objects, adding motion to existing in-the-wild 3D assets, and synthesis of motion across unbounded, large-scale scenes. Finally, designing automated text-to-4D metrics by temporally extending text-to-3D metrics [96] is an important direction for future work.
@@
Our approach can automatically generate realistic, animated 3D scenes; such techniques can be misused to promulgate realistic fake content for the purposes of misinformation. We condemn misuse of generative models in this fashion and emphasize the importance of research to thwart such efforts (see, e.g., Masood et al. [55] for an extended discussion of this topic).
Acknowledgements